OpenStudy (ashleyy_smith):

Use substitution to solve the system. x = 3y - 17 −5x + 2y = 33

OpenStudy (photon336):

you see something here? you have an expression for x already x = 3y-17 what do you think you can do with that?

rebeccaxhawaii (rebeccaxhawaii):

put both in standard form first

rishavraj (rishavraj):

plug the value of x u got in 1st equation in the 2nd equation

rebeccaxhawaii (rebeccaxhawaii):

looks like @Photon336 got the question first ;'c

OpenStudy (ashleyy_smith):

woah so many people haha :o

OpenStudy (ashleyy_smith):

divide right?

rishavraj (rishavraj):

no no..... u jst need to plug the value in 2nd equation

OpenStudy (ashleyy_smith):

how :s

rishavraj (rishavraj):

like in the first equation u got x = 3y - 17 now just plug tht in 2nd equation

rishavraj (rishavraj):

@Ashleyy_Smith http://www.purplemath.com/modules/systlin4.htm

rishavraj (rishavraj):

so on substituting the value of x in 2nd equation.... it would be \[-5(3y -17) + 2y = 33\]

OpenStudy (ashleyy_smith):

Do I distribute now?

rishavraj (rishavraj):

now u got the equation with only one variable..... deal with it... :P get y .... plug its value in any of the equation and get x

rebeccaxhawaii (rebeccaxhawaii):

so what @rishavraj is showing you is the substitution way there is also a elimination way

OpenStudy (ashleyy_smith):

yes but I have to give the answer in substitution form :s

rebeccaxhawaii (rebeccaxhawaii):

its a method

rishavraj (rishavraj):

answer is gonna remain the same -_- @Ashleyy_Smith bt if u supposd to show ur work ...then u neeed to solve it by substitution method

OpenStudy (ashleyy_smith):

so -15+85+2=33 right?

rishavraj (rishavraj):

wheres y in the equation ??? -_-

rishavraj (rishavraj):

-5(3y - 17)y + 2y = 33 -15y + 85 + 2y = 33 solve for y

OpenStudy (ashleyy_smith):

oops after the 15 and the 2

rishavraj (rishavraj):

okkk go on solve it ...

OpenStudy (ashleyy_smith):

I add the y's first right?

rishavraj (rishavraj):

go on....... just get me the y

OpenStudy (ashleyy_smith):

-13y

rishavraj (rishavraj):

so its -13y + 85 = 33 u mean tht??

OpenStudy (ashleyy_smith):

yes

rishavraj (rishavraj):

go on...... then get y

OpenStudy (ashleyy_smith):

I got a long decimal when I divided 72 by 33

rishavraj (rishavraj):

why decimal..... how do u solve y + 2 = 5 ????

rishavraj (rishavraj):

ok tell me whts the value of y , when y + 2 = 5 ??

OpenStudy (ashleyy_smith):

would it be 5?

rishavraj (rishavraj):

for y + 2 = 5 ????y = 5>??

OpenStudy (ashleyy_smith):

ohhhhh oops my bad 3

rishavraj (rishavraj):

see u need to isolate the variable on one side of the sign of equality..... so like y + 2 = 5 ,...... u need to subtract 2 on both sides... so u will be having y + 2 - 2 = 5 -2 y = 3 u get tht ?? @Ashleyy_Smith

rishavraj (rishavraj):

ooo good..... now try -13y + 85 = 33

rishavraj (rishavraj):

wht do u think would be the first step ??

OpenStudy (ashleyy_smith):

oops

rishavraj (rishavraj):

hmmmm ??

rishavraj (rishavraj):

solve -13y + 85 = 33

OpenStudy (ashleyy_smith):

uuummm. would I subtract 33 from 85?

rishavraj (rishavraj):

as I told u need to isolate the variable .... so u need to subtract __ on both sides

OpenStudy (ashleyy_smith):

so 72 and 20?

rishavraj (rishavraj):

wht is tht now??? -_- u simply need to subtract 85 on both sides

OpenStudy (ashleyy_smith):

oh

rishavraj (rishavraj):

so it would be -13y = -52 now wht??

OpenStudy (ashleyy_smith):

divide and get -4

rishavraj (rishavraj):

means -y = -4

rishavraj (rishavraj):

so whts the value of y??

OpenStudy (ashleyy_smith):

u just said y = -4 :o

rishavraj (rishavraj):

its -y = -4

OpenStudy (ashleyy_smith):

so u mean of x?

OpenStudy (ashleyy_smith):

oh

rishavraj (rishavraj):

theres a negative sign -y = -4 so y = ??

OpenStudy (ashleyy_smith):

4

rishavraj (rishavraj):

good ...now plug the value in any equation....lol prefer the 1st one in this case

OpenStudy (ashleyy_smith):

x=3y-17 x= 3(4) - 17 x = 12y - 17

OpenStudy (ashleyy_smith):

right?

rishavraj (rishavraj):

x = 12 -17 -_- so x = ??

OpenStudy (ashleyy_smith):

5

OpenStudy (ashleyy_smith):

-5

rishavraj (rishavraj):

oooo coolll... so its x = -5 and y = 4

rishavraj (rishavraj):

good work :))

rishavraj (rishavraj):

@Ashleyy_Smith u got this Substitution method??

OpenStudy (ashleyy_smith):

yes :) ill practice more on it tomorrow

rishavraj (rishavraj):

good :))) All the Best

OpenStudy (ashleyy_smith):

thank u!